OpenStudy (anonymous):

simplify ^4 sqrt 27 (times) ^4 sqrt 3

OpenStudy (anonymous):

\[\sqrt[4]{27*\sqrt[4]{3}} = \sqrt[4]{3^3*3^{1/4}} = \sqrt[4]{3^{13/4}}\]\[3^{13/16}\]